MILLIMETER- WAVE SPECTRUM OF PYRIDAZINE

\underline{BRENT K. AMBERGER}, BRIAN J. ESSELMAN, JOSH D. SHUTTER, R. CLAUDE WOODS and ROBERT J. McMAHON, The University of Wisconsin - Madison Department of Chemistry, 1101 University Avenue, Madison, WI 53705.

Pyridazine, the ortho dinitrogen substituted analogue of benzene, was chosen as an inaugural project for our newly modernized millimeter-wave absorption spectrometer. Pyridazine has a substantial dipole of 4.22 debye, and is thus a candidate for astronomical detection. If detected, it would be among the first 6π electron aromatic systems to be observed in space. Spectra in the range of 260-360 GHz were collected at room temperature. Pure rotational spectra for 10 vibrational states and 4 naturally occurring isotopomers are assigned and reported. Among the vibrationally excited states, both Coriolis and Fermi coupling is observed. Anharmonic vibration-rotation calculations using the CFOUR package are used to find and assign transitions for the vibrationally excited states. These calculations at the CCSD(T)/cc-pVTZ level provide excellent predictions for the rotational constants of these excited states.